# Foundation Models for the Brain and Body

### v1

@@ Abstract
Subject-specific distribution shifts represent an important obstacle to the development of foundation models for EEG decoding. To address this, we propose Subject-Conditioned Layer, an adaptive layer designed as a drop-in replacement for standard linear or convolutional layers in any neural network architecture. Our layer captures subject-specific variability by decomposing its weights into a shared, subject-invariant component and a lightweight, low-rank correction unique to each subject. This explicit separation of general knowledge from personalized adaptation allows existing models to become robust to subject shifts. Empirically, models equipped with our layer outperform both a shared-weight-only model (subject-agnostic model) and the average of individually trained subject-specific models. Consequently, the Subject-Conditioned Layer offers a practical and scalable path towards building effective cross-subject foundation models for EEG.
@@ 1 Introduction
Developing a foundation model for decoding complex brain signals like Electroencephalography (EEG) is a central challenge in neuroscience and machine learning [2]. Progress is slowed by the inherent properties of EEG data: low signal-to-noise ratio, non-stationarity, and, most critically, high inter-subject variability [35]. Neural patterns vary so widely across individuals, and even within the same individual over time, that data is effectively fragmented, exacerbating the problem of data scarcity [3, 25]. Consequently, large models trained on pooled data from many subjects often fail to outperform simpler models tailored to individuals, stalling progress towards generalizable solutions.
@@
This situation contrasts sharply with fields like computer vision (CV) and natural language processing (NLP). While these fields have also relied on domain-specific architectures, the Transformer [34] provided a universal paradigm that unlocked predictable scaling laws, enabling massive foundation models to emerge [13, 37]. EEG decoding has yet to benefit from similar scaling. The challenge of cross-subject generalization has kept effective dataset sizes small, preventing the successful application of large-scale, general-purpose architectures. As a result, the field remains dominated by specialized models that have not yet converged to a standard, foundational approach [2, 12].
@@
To bridge this gap, we address the problem of subject-specific distribution shifts, which we identify as the primary obstacle to building foundation models for multi-subject problems such as EEG classification. We propose Subject-Conditioned Layer, an approach that decomposes the parameters of a neural network into two components: a set of base weights shared across all subjects and a subject-specific, low-rank correction that efficiently captures individual variability. The Subject-Conditioned Layer module is designed as a drop-in replacement for standard linear or convolutional layers, making it universally applicable to a wide range of model architectures with applications in EEG decoding and beyond.
Our contributions are threefold: We introduce Subject-Conditioned Layer, a novel adaptive layer that explicitly separates shared knowledge from subject-specific signatures and is designed as a drop-in replacement for standard linear and convolutional layers in any neural network architecture. We provide qualitative evidence through embedding visualization that our layer successfully disentangles subject-invariant neural patterns from individual subject signatures, confirming the mechanism behind our method’s success. Through a comprehensive evaluation on the BCI Competition IV datasets, we demonstrate that both CNN and ViT like architectures equipped with Subject-Conditioned Layer outperform strong baselines: (i) a subject-agnostic model, (ii) the average of subject-specific models, and (iii) the average of subject-specific LoRA models. The code is publicly available  [^1] https://github.com/timonkl/SubjectConditionedLayer.
Figure 1: Latent Representations Colored by Subject ID vs. Target Labels. The top row compares t-SNE projections of representations derived from (Left) the shared weights ($W_{\text{general}}$), (Middle) the subject-specific weights ($W_{s}$), and (Right) the fused representations from the complete Subject-Conditioned Layer. In this row, all points are colored by subject ID. The bottom row presents the same data points, but colored according to their ground-truth target labels. This comparison clearly shows that the sub-clusters formed by our full model (top right) correspond directly to the target classes (bottom right), demonstrating that the model learns discriminative, task-relevant features within each subject’s representation space. All representations are obtained from the classification token predicted by the Patched Brain Transformer, with a linear classification head used for visualization clarity.
@@ 2 Related Work
Our proposed method addresses the challenge of inter-subject variability in EEG decoding by learning both shared and subject-specific representations. This approach is closely related to established paradigms nonlinear mixed-effects (NLME) modeling and in machine learning, primarily Multi-Task Learning (MTL) and Mixture of Experts (MoE).
Nonlinear Mixed-Effects Modeling. Our approach can be viewed through the lens of nonlinear mixed-effects (NLME) models [21, 22]. In this framing, the shared network $W_{\text{general}}$ learns the fixed, population-level effects, while the subject-specific low-rank adapters $W_{s}$ capture the random effects, parsimonious deviations for each individual. The low-rank constraint on the adapters is analogous to the low-dimensional random effects assumption in classical NLME, which encourages partial pooling of information and borrows statistical strength across subjects. This perspective provides a principled justification for using LoRA [14] as a parameter-efficient method to model subject-specific variations, which is crucial when per-subject data is limited.
Multi-Task Learning. In MTL, a model is trained on several related tasks simultaneously, with the goal of improving performance by leveraging shared information across tasks [5, 23, 9]. Viewing the decoding for each subject as a distinct task, our model fits directly into the MTL framework. It employs a form of hard parameter sharing, where the general weight matrix, $W_{\text{general}}$, learns a common feature space shared across all subjects. The subject-specific weights, $W_{s}$, act as task-specific layers that adapt this shared representation to the unique neural patterns of each individual. This architecture allows the model to learn robust, generalizable features from the entire population while still retaining the flexibility to specialize for each subject [38].
Mixture of Experts. MoE models utilize multiple expert networks and a gating mechanism that determines which expert processes a given input [15]. Sparsely-gated MoE models, in particular, activate a subset of experts for each input, enabling the scaling of model capacity without a proportional increase in computation [31, 10]. Our subject-aware layer can be interpreted as a specialized, non-learned MoE. Here, the subject identity serves as the gating signal, deterministically routing the input to its corresponding expert, the pathway defined by $W_{s}$. Unlike traditional MoE, our gating is not a learned function but is provided by the input via the mask M. This simplifies the model by removing the need to learn a routing policy and directly addresses the known source of variation (i.e., subject identity).
Dynamical low-rank. Since our method explicitly models subject-specific variability as lightweight low-rank corrections, it naturally connects to the broader line of dynamical low-rank research [26, 28, 29]. An exciting direction for future work is to explore whether subject adaptation in EEG can benefit from recent advances in geometry-aware training [36, 30], and robust compression strategies [29], thereby unifying personalization with efficient large-scale deployment.
@@ 3 On the Difficulty of Multi-Subject Training
We consider a classification problem where data is collected from multiple subjects. Since the signal contains subject-specific characteristics, our goal is to learn a general model that captures class-discriminative structures across subjects. Let $\mathcal{X}_{s}=\{x_{sj}\}_{j=1}^{M_{s}}$ and $\mathcal{Y}_{s}=\{y_{sj}\}_{j=1}^{M_{s}}$ contain subject-specific features and labels, respectively, for subjects $s\in\{1,\dots,N\}$. In short, we write $\mathcal{D}=(\mathcal{X},\mathcal{Y})=\bigcup_{s=1}^{N}\mathcal{D}_{s}=\bigcup_{s=1}^{N}(\mathcal{X}_{s},\mathcal{Y}_{s})$.
@@
which states that once the latent representation $\mathcal{Z}$ is known, the participant index $s$ and the datapoint $x_{sj}$ provide no additional information about the label $y_{sj}$. This conditional independence assumption implies that $\mathcal{Z}$ serves as a sufficient statistic of $x_{sj}$ to predict $y_{sj}$. A key step is to verify whether this property holds for the proposed method.
@@ 4 Method
Figure 2: Subject-Conditioned Layer. The input $X$ is processed by a shared general weight matrix $W_{\text{general}}$. For each subject, a subject-specific low-rank adapter, parameterized by matrices $A_{s}$ and $B_{s}$ of rank $r$, computes a correction that is added to the output of the general weight matrix, yielding the layer output $\bar{X}$.
4.1 Subject-Conditioned Layer
@@
where $W_{\text{general}}\in\mathbb{R}^{m\times n}$ and $W_{s}\in\mathbb{R}^{m\times n}$ are shared and subject-specific weight matrices, respectively. The mask matrices $M_{s}$ with binary entries extract subject-$s$-specific samples from the batch data $X$.
Note that this approach may lead to a reduction in batch sizes for subject-dependent weights $W_{s}$, if the number of subjects $N$ increases.
As a remedy, adapting the total batch size was proposed in [31].
@@ 4.2 Low Rank approximation for Subject-Dependent Weights
Low-rank adaptation (LoRA) [14] is a well-established method for fine-tuning models in few-shot settings. Compared to full-rank training, LoRA shows superior performance on limited data [27], while also substantially reducing the number of training parameters and the associated memory consumption. We define subject-specific weights as low-rank matrix decompositions
@@
with $A_{s}\in\mathbb{R}^{n\times r},\;B_{s}\in\mathbb{R}^{r\times m}$ and rank $r$. Similarly to [14], we initialize $A_{s}=\mathcal{N}(0,\sigma^{2})$, and $B_{s}=0$ for all subjects $s$.
4.3 Regularization of Subject-Dependent Weights
A central challenge is to ensure that the general weight matrix $W_{\text{general}}$ learns shared patterns while the subject-specific weights $W_{s}$ only capture individual deviations. To prevent $W_{s}$ from learning general representations, we employ a regularization strategy. First, we constrain the complexity of the subject-specific corrections by enforcing a relatively small rank $r$, as defined in Equation 5. See Tables 2 and 3 for exact values.
Second, we introduce an optimization-based regularization method by scaling $W_{s}$ with $\frac{\alpha}{r}$, where $\alpha$ is a hyperparameter. This scaling adjusts the effective learning rate of the adapter matrices relative to the general weights. For $\alpha<r$, the adapters are updated conservatively, encouraging reliance on $W_{\text{general}}$. In contrast to $\alpha>r$, where the adapters are emphasized.

### v2

# Mitigating Subject Dependency in EEG Decoding with Subject-Specific Low-Rank Adapters

@@ Abstract
Subject-specific distribution shifts represent a fundamental obstacle to developing foundation models for brain decoding. We propose the Subject-Specific Low-Rank Adapter (SuLoRA), a drop-in replacement for standard linear or convolutional layers that captures inter-subject variability by decomposing weights into a shared, subject-invariant component and a lightweight, low-rank correction unique to each subject. This explicit separation enables existing architectures to become robust to subject shifts without architectural redesign. We evaluate SuLoRA on MEG speech perception and EEG motor imagery tasks across CNN and transformer architectures. In the speech decoding task, SuLoRA exceeds the baseline performance with half of the parameters. On motor imagery dataset, SuLoRA outperforms both subject-agnostic models and independently trained subject-specific models. SuLoRA offers a practical path towards effective cross-subject foundation models for brain signal applications.
Machine Learning, ICML
@@ 1 Introduction
Developing a foundation model for decoding complex brain signals such as Electroencephalography (EEG) and magnetoencephalography (MEG) is a central challenge in neuroscience and machine learning (Aristimunha et al., 2025c). Progress is slowed by the inherent properties of EEG data: low signal-to-noise ratio, non-stationarity, and, most critically, high inter-subject variability (Wei et al., 2022). Neural patterns vary so widely between individuals and even within the same individual over time. The data are effectively fragmented, amplifying the problem of data scarcity (Banville et al., 2025; Bomatter and Gouk, 2025; Schmid et al., 2024). Consequently, large models trained on pooled data from many subjects often fail to outperform simpler models tailored to individuals, slowing progress towards generalizable solutions.
@@
This situation contrasts sharply with fields of natural language processing or computer vision. Although these fields have also relied on domain-specific architectures, the Transformer (Vaswani et al., 2023) provided a universal paradigm that unlocked predictable scaling laws, allowing massive foundation models to emerge (Hoffmann et al., 2022; Zhai et al., 2022). Brain decoding has yet to benefit from similar scaling. The challenge of cross-subject generalization has kept effective dataset sizes small, preventing the successful application of large-scale, general-purpose architectures. As a result, the field remains dominated by specialized models that have not yet converged to a standard, foundational approach (Aristimunha et al., 2025c; Guetschel et al., 2024).
@@
To bridge this gap, we address the problem of subject-specific distribution shifts, which we identify as the primary obstacle to building foundation models for multi-subject brain-signal applications. We propose the Subject-Specific Low-Rank Adapter (SuLoRA), an approach that decomposes the parameters of a neural network into two subsets: a set of base weights shared across all subjects and a subject-specific, low-rank correction that efficiently captures individual variability. The SuLoRA module is designed as a drop-in replacement for standard linear or convolutional layers, making it universally applicable to a wide range of model architectures with applications in EEG/MEG decoding and beyond.
Our contributions read as follows:
  • We introduce SuLoRA, a novel adaptive layer that explicitly separates shared knowledge from subject-specific signatures, designed as a drop-in replacement for linear and convolutional layers in any architecture.
  • We provide qualitative evidence through embedding visualization that SuLoRA successfully disentangles subject-invariant neural patterns from individual signatures.
  • We demonstrate through comprehensive evaluation on speech perception (MEG) and motor imagery (EEG) tasks that models equipped with SuLoRA outperform: (i) subject-agnostic models, (ii) independently trained subject-specific models, and (iii) independent subject-specific LoRA models.
Code [^1] https://github.com/timonkl/SubjectConditionedLayer and data are publicly available.
@@ 2 Related Work
Scaling laws in natural language processing and computer vision have established a reliable paradigm for improving predictive performance by jointly increasing model size, training compute, and dataset scale (Kaplan et al., 2020; Hoffmann et al., 2022; Zhai et al., 2022). In contrast, recent work on scaling laws for neural decoding shows that performance improves with increased training data per subject, but does not scale proportionally with the number of subjects (Banville et al., 2025). This presents a fundamental obstacle to developing large-scale, multi-subject foundation models for brain decoding, where aggregating data across individuals is both necessary and desirable. Despite these challenges, Leave-One-Subject-Out (LOSO) benchmarks (Wang et al., 2020; Kwon et al., 2020; Zhang et al., 2025) demonstrate that cross-subject transfer is generally possible, although with notable performance degradation. Taken together, these findings suggest that brain decoding representations comprise both transferable subject-agnostic components and subject-specific components that limit performance gains when scaling across subjects.
To address inter-subject variability in brain decoding, a wide range of approaches has been proposed. Early methods introduce subject-dependent bias terms or linear transformations applied to learned features (Ma et al., 2019). Recent studies investigate inter-subject variability via explicit adaptation mechanisms, such as subject-conditioned batch normalization and adaptive feature modulation. In particular, (Liang et al., 2024) learn a shared subject-invariant feature space while applying lightweight subject-specific transformations to correct individual distortions, enabling cross-subject decoding using a small amount of target-subject data.
More recent work in speech decoding (lévy2025braintotextdecodingnoninvasiveapproach; d’Ascoli et al., 2024; Shapovalenko and Auster, 2025) introduces a subject-specific linear layer applied early in the network. This layer performs a channel-wise transformation that aligns subject-specific representations into a shared latent space while keeping the remaining parameters shared across subjects. However, this approach introduces a relatively large number of additional parameters that grow linearly with the number of subjects. Ablating these subject-specific layers leads to significant accuracy drops (Défossez et al., 2023b), underscoring the importance of explicitly modeling subject-dependent effects.
Our SuLoRA directly addresses inter-subject variability by explicitly learning both shared and subject-specific representations. It introduces a set of global parameters augmented with lightweight, subject-specific corrections. This design is inspired by nonlinear mixed-effects models (Lindstrom and Bates, 1990; Pinheiro and Bates, 2000), which jointly model population-level effects and parsimonious individual-level deviations via random effects. The approach also naturally aligns with multi-task learning (MTL) (Caruana, 1997; Evgeniou and Pontil, 2004; Ruder, 2017), where shared structure across related tasks is leveraged while allowing task-specific variation.
Conceptually, our approach is related to Mixture-of-Experts (MoE) models (Jiang et al., 2024a; Dai et al., 2024; Riquelme et al., 2021), which decompose computation across multiple expert networks and use a gating mechanism to select experts for each input (Jacobs et al., 1991). Sparsely gated MoE variants enable scaling model capacity without proportional increases in computational cost by activating only a subset of experts per example (Shazeer et al., 2017; Fedus et al., 2022). MoE architectures have been adapted for brain decoding: EEGMamba (Gui et al., 2024) employs a task-aware learned gating network together with a universal expert to capture shared EEG structure, while MoGE (Liu et al., 2024) and VMoGE (Ding et al., 2025) explore graph-based experts that assign channels or frequency bands to specialized experts. In brain decoding, subject identity provides a deterministic and known source of variation, acting as a fixed gating signal that routes each input to its corresponding expert pathway. Unlike conventional MoE models that rely on learned routing policies, our approach exploits this prior structure directly, simplifying the architecture while explicitly modeling the source of inter-subject variability that limits cross-subject scaling.
Figure 1: SuLoRA architecture. The input $X$ is processed by a shared weight matrix $W_{\text{general}}\in\mathbb{R}^{m\times n}$ capturing subject-invariant patterns. Each subject s has a dedicated low-rank adapter, parameterized by matrices $A_{s}\in\mathbb{R}^{n\times r}$ and $B_{s}\in\mathbb{R}^{r\times m}$, which computes a subject-specific correction added to the shared output. Unlike standard MoE models, SuLoRA uses known subject identity to deterministically select the appropriate adapter at both training and inference.
Our implementation leverages a shared full-rank expert that is always active, together with subject-specific LoRA modules. LoRA (Hu et al., 2021), originally introduced as a parameter-efficient fine-tuning method that adds low-rank updates to pretrained weights, has been successfully applied across modalities and tasks, offering strong empirical performance with minimal memory and compute overhead. Prior work has explored design choices such as rank selection (Zangrando et al., 2024; schotthöfer2024federateddynamicallowranktraining) and robust compression strategies (schotthöfer2025dynamicallowrankcompressionneural). In our framework, LoRA serves as the mechanism for implementing subject-specific corrections, enabling fine-grained adaptation to individual subjects while tightly controlling the number of additional parameters. It has also been shown that LoRA enables parameter-efficient fine-tuning of large brainwave models, achieving competitive EEG decoding performance with minimal trainable parameters (Lee et al., 2025).
@@ 3 On the Difficulty of Multi-Subject Training
In this section, we focus solely on a classification problem in which data is collected from multiple subjects. Since the signal contains subject-specific characteristics, our goal is to learn a general model that captures class-discriminative structures across subjects. Let $\mathcal{X}_{s}=\{x_{sj}\}_{j=1}^{M_{s}}$ and $\mathcal{Y}_{s}=\{y_{sj}\}_{j=1}^{M_{s}}$ contain subject-specific features and labels, respectively, for subjects $s\in\{1,\dots,N\}$. In short, we write $\mathcal{D}=(\mathcal{X},\mathcal{Y})=\bigcup_{s=1}^{N}\mathcal{D}_{s}=\bigcup_{s=1}^{N}(\mathcal{X}_{s},\mathcal{Y}_{s})$.
@@
which states that once the latent representation $\mathcal{Z}$ is known, the participant index $s$ and the datapoint $x_{sj}$ provide no additional information about the label $y_{sj}$. This conditional independence assumption implies that $\mathcal{Z}$ serves as a sufficient statistic of $x_{sj}$ to predict $y_{sj}$.
@@ 4 Method
4.1 Subject-Specific Low-Rank Adapter
@@
where $W_{\text{general}}\in\mathbb{R}^{m\times n}$ and $W_{s}\in\mathbb{R}^{m\times n}$ are shared and subject-specific weight matrices, respectively. The mask matrices $M_{s}$ with binary entries extract subject-$s$-specific samples from the batch data $X$. Note that this approach may lead to a reduction in batch sizes for subject-dependent weights $W_{s}$, if the number of subjects $N$ increases. As a remedy, adapting the total batch size was proposed in (Shazeer et al., 2017).
@@ 4.2 Low Rank approximation for Subject-Dependent Weights
Low-rank adaptation (LoRA) (Hu et al., 2021) is a well-established method for fine-tuning models in few-shot settings. Compared to full-rank training, LoRA shows superior performance on limited data (Schotthöfer et al., 2025), while also substantially reducing the number of training parameters and the associated memory consumption. We define subject-specific weights as low-rank matrix decompositions
@@
with $A_{s}\in\mathbb{R}^{n\times r},\;B_{s}\in\mathbb{R}^{r\times m}$ and rank $r$. Unlike standard LoRA, we initialize $A_{s}\sim\mathcal{N}\left(0,\frac{2}{r}\right)$ and $B_{s}\sim\mathcal{N}(0,0.01^{2})$ for all subjects $s$. Initializing $B_{s}=0$, as in standard LoRA, led to reduced performance, likely due to vanishing gradients through $A_{s}$ in early training.
4.3 Scaling of Subject-Specific Adapters
Similar to the original LoRA, we introduce a scaling of $W_{s}$ with $\frac{\alpha}{r}$, where $\alpha$ is a hyperparameter. This scaling adjusts the effective learning rate of the adapter matrices relative to the general weights. For $\alpha<r$, the adapters are updated conservatively, encouraging reliance on $W_{\text{general}}$, in contrast to $\alpha>r$, where the adapters are emphasized.
4.4 Regularization of Subject-Specific Adapters
A central challenge is to ensure that the general weight matrix $W_{\text{general}}$ learns shared patterns while the subject-specific weights $W_{s}$ only capture individual deviations. To prevent $W_{s}$ from learning general representations, we employ a regularization strategy. We constrain the complexity of the subject-specific corrections by enforcing a relatively small rank $r$, as defined in Equation 5. See Tables 4 and 5 for exact values.